A penny falls from the top of a building, and hits the ground going 30 m/s. How tall is the building? Ignore

air resistance.


1
Expert's answer
2021-09-20T17:40:56-0400

Initial Velocity = u = 0 m/s

Final Velocity = v = 30 m/sec (at the time of hit the ground)

g = 9.8 m/sec2


3rd equation of motion:

"v^2 = u^2 + 2gh"

30 * 30 = 0 + 2*9.8 * h

"h = \frac{30*30}{2*9.8} = 45.91 meters"


Final Answer: The building is 45.9 meters tall.